OpenStudy (anonymous):

NEED HELP ASAP. WILL FAN AND MEDAL. A firecracker shoots up from a hill 150 feet high, with an initial speed of 110 feet per second. Using the formula H(t) = -16t^2 + vt + s, approximately how long will it take the firecracker to hit the ground? Eight seconds Nine seconds 10 seconds 11 seconds

OpenStudy (anonymous):

@pooja195 @Concentrationalizing

OpenStudy (anonymous):

If the hill is 150 feet high, then that's essentially saying that when t = 0, h(t) should equal 150. An initial speed of 110 feet just tells you that v = 110. Using that information, you can rewrite h(t) as a quadratic in t only in which then you would just factor and solve for t.

OpenStudy (anonymous):

Ok, can you walk me through it? :)

OpenStudy (anonymous):

Well, if you plug in t = 0, what do you get?

OpenStudy (anonymous):

h(0) = -16 + s?

OpenStudy (anonymous):

Well, if t is 0, the -16 goes away also. \(-16(0)^{2} +v(0) + s = s\) . In the end, the problem just tells you that s = 150 and v = 110.

OpenStudy (anonymous):

Ok :) so what next?

OpenStudy (anonymous):

Well, that means you have \(h(t) = -16t^{2} +110t + 150\) So, h(t) represents height. You plug in a time value and you get a height for an answer. The problem wants to know when you hit the ground. Well, if you hit the ground, then the height is 0. So we want to know what time value will make h(t) equal 0. Following me so far?

OpenStudy (anonymous):

Oh ok :D I think so.

OpenStudy (anonymous):

Mhm. So if h(t) needs to be 0, then you're solving this: \(-16t^{2} + 110t + 150 = 0\)

OpenStudy (anonymous):

Ok just give me a minute...

OpenStudy (anonymous):

Alrighty

OpenStudy (anonymous):

Ok, I think I got lost. I only got to the beginning: -16t^2 + 110t = -150. :(

OpenStudy (anonymous):

After that I don't know what to do.

OpenStudy (anonymous):

Well, this is a quadratic equation, so it needs to be factored.

OpenStudy (anonymous):

Ok, so -2t(8t - 55) = -150?

OpenStudy (anonymous):

Nope. Have you seen before how an equation like \(x^{2} + 5x + 6\) turns into \((x+2)(x+3)\)?

OpenStudy (anonymous):

Yes, I have. :) So I need to do that here?

OpenStudy (anonymous):

Yeah, its that idea. So itll factor and look like ( )( ) = 0 Now it's just knowing how to factor in that way.

OpenStudy (anonymous):

Hmm. So would it start to look like (-4t +/- ___)(-4t +/- ___)?

OpenStudy (anonymous):

Actually, this doesnt factor cleanly. Since the question said "approximately", we won't be able to factor in the way I was thinking. We're actually forced to quadratic formula on this, I apologize.

OpenStudy (anonymous):

Oh ok! That's fine :)

OpenStudy (anonymous):

Youre okay with doing quadratic formula correctly? :)

OpenStudy (anonymous):

Yes. Here give me just a few minutes ;) things around here are getting a bit crazy XD

OpenStudy (anonymous):

Ok so is it -8.04 or 1.17?? Oh so then it would be eight seconds??

OpenStudy (anonymous):

Sorry I took so long lol

OpenStudy (anonymous):

I didnt do the full math, but itd be 8 seconds, yeah :3

OpenStudy (anonymous):

Ok thanks! I'll give you a medal but I'm already a fan ;)

OpenStudy (anonymous):

Lol, okay. As long as it makes sense.

OpenStudy (anonymous):

Yes it does now ;) because of you. Thanks so much! Ttyl :)

OpenStudy (anonymous):

Good luck :3